# PCF Learned Sort: a Learning Augmented Sort Algorithm with $\mathcal{O}(n\log\log n)$ Expected Complexity

### v1

@@ Appendix A Proofs
$Y_{jr}=\begin{cases}1&(j\in\mathcal{S}_{r})\\ 0&(\mathrm{else})\end{cases},~{}~{}~{}~{}~{}~{}~{}~{}Z_{jr}=\begin{cases}1&(j\in\mathcal{T}_{r})\\ 0&(\mathrm{else})\end{cases},$ (28)
$Y_{r}=\sum_{j=1}^{\alpha}Y_{jr},~{}~{}~{}~{}~{}~{}~{}~{}Z_{r}=\sum_{j=1}^{\alpha}Z_{jr}.$ (29)
If the size of the bucket to which $e_{r}$ is allocated is greater than or equal to $\delta$, then the following holds:
@@
Using Lemma A.1, we can prove Lemma 3.4.
@@
[…]
@@
$10\leq 1-\frac{\alpha}{\gamma\mathbb{E}[Y_{r}]}<$ 1 . (37)
@@
Thus, by defining $E_{r}$ to be the event “$e_{r}$ is allocated to a bucket with size greater than or equal to $\delta$,” from Lemma A.1,

### v3

@@ Appendix A Proofs
$Y_{jr}=\begin{cases}1&(j\in\mathcal{S}_{r})\\ 0&(\mathrm{else})\end{cases},~~~~~Z_{jr}=\begin{cases}1&(j\in\mathcal{T}_{r})\\ 0&(\mathrm{else})\end{cases},~~~~~Y_{r}=\sum_{j=1}^{\alpha}Y_{jr},~~~~~Z_{r}=\sum_{j=1}^{\alpha}Z_{jr}$ (18)
When using $\mathcal{M}_{\mathrm{PCF}}$, if the size of the bucket to which $e_{r}$ is allocated is greater than or equal to $\delta$, the following holds:
@@
Using Lemma˜A.1, we can prove Lemma˜3.4.
@@
Let $q=\max\limits_{y}\int_{y}^{y+\Delta}f_{\mathcal{I}}(x)dx$ (where $y$ is a value such that $(y,y+\Delta)\subseteq\mathcal{I}$). Then, from $\sigma_{1}\leq f(x)\leq\sigma_{2}$ for all $x\in\mathcal{I}$,
$\begin{split}q&\leq\frac{\max_{y}\int_{y}^{y+\Delta}f(y)dy}{\int_{\mathcal{I}}f(x)dx}\\ &\leq\frac{\max_{y}\int_{y}^{y+\Delta}\sigma_{2}dy}{\int_{\mathcal{I}}\sigma_{1}dx}\\ &\leq\frac{\sigma_{2}\Delta}{\sigma_{1}(x_{\mathrm{max}}-x_{\mathrm{min}})}\\ &=\frac{\sigma_{2}}{\sigma_{1}\beta}.\end{split}$ (23)
@@
$0\leq 1-\frac{\alpha}{\gamma\mathbb{E}[Y_{r}]}<1.$ (26)
@@
Thus, by defining $E_{r}$ to be the event $e_{r}$ is allocated to a bucket with size greater than or equal to $\delta$,” from Lemma˜A.1,